OpenStudy (anonymous):

Please help me to Solve this problem. Susan purchased some municipal bonds yielding 7% annually and some certificates of deposit yielding 9% annually. If Susan's investment amounts to $19,000 and the annual income is $1590, how much money is invested in bonds and how much is invested in certificates of deposit? Seleccione una respuesta. a. $13,000 in bonds; $6000 in certificates of deposit b. $5500 in bonds; $13,500 in certificates of deposit c. $13,500 in bonds; $5500 in certificates of deposit Incorrecto d. $6000 in bonds; $13,000 in certificates of deposit

OpenStudy (anonymous):

Let bonds = x and certificates = y, then x + y = 19,000 and .07x +.09y = 1590 We now have a system of equations and can solve by substitution. x = 19000 - y If we sub x into the bottom equation, we get .07(19000 - y) + .09y = 1590 1330 - .07y + .09y = 1590 .02y = 260 y = 13,000 in certificates Now sub y into the top equation and we get x = 19000 - 13000 x = 6000 in bonds

OpenStudy (anonymous):

Thanks